(6/y+4)+1=(5/2y+8)
I thought that the next step would look like:
6(2y+8)+(y+4)(2y+8)=5(y+4) but the answer in the back of the book does not
correspond with this once I finish solving it.
The final answer should be -15/2.

Answer by Alan3354(69443)

6/y+4)+1=(5/2y+8)
I thought that the next step would look like:
6(2y+8)+(y+4)(2y+8)=5(y+4) but the answer in the back of the book does not
correspond with this once I finish solving it.
The final answer should be -15/2.
----------------
The DEN of the right side is 2x the one on the left.
-------------
6/y+4)+1=(5/2y+8)
Multiply all terms by the right side DEN 2y+8
12 + 2y+8 = 5
2y + 20 = 5
2y = -15
y = -15/2